OpenStudy (anonymous):

Simplify f(x)=(3x^2-4x+1)/(3x-1)Using complete sentences, explain why f(1) = 0, f(0) = –1, and f(–1)=–2, yet f( 1/3 ) is undefined. Make sure to show your work.

OpenStudy (anonymous):

@myko ??? Can you help??

OpenStudy (anonymous):

\(\frac{(3x^2-4x+1)}{(3x-1)}=\frac{(3x-1)(x-1)}{(3x-1)}=x-1\) so : f(1) = 0, f(0) = –1, and f(–1)=–2 and f( 1/3 ) is not alowed , because it makes denominator in the original expretion 0

OpenStudy (anonymous):

Okay so because it's 0 it's undefined? What should I say about why the f(1)=0 and f(0)=-1?

OpenStudy (anonymous):

Does that also mean it's a restriction?

OpenStudy (anonymous):

you function f(x)=(3x^2-4x+1)/(3x-1) was simplified to f(x)=x-1. So: f(1)=1-1=0 f(0)=0-1=-1 and x=1/3 is not alowed, so it's a restriction

OpenStudy (anonymous):

THANKS @myko !!!!! :D

OpenStudy (skullpatrol):

Do you understand why division by 0 is undefined?

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

When it equals 0 it means it's undefined

OpenStudy (skullpatrol):

Yes that is true, but do you know why?

OpenStudy (anonymous):

No why?

OpenStudy (skullpatrol):

In math when we say something is undefined we mean that it is not part of the definition. In the case of division using 0 as a denominator is strictly not allowed.

OpenStudy (anonymous):

if \(\huge \frac{a}{0}=b\) then it should be \(\huge b*0=a\). But here is no number \(\huge b\) that multiplied by \(\huge 0\) gives \(\huge a\), because any number multiplied by 0 equals 0

OpenStudy (anonymous):

@skullpatrol

OpenStudy (anonymous):

that's why

OpenStudy (anonymous):

Ok gothcha! Thanks guys!! :) :)

OpenStudy (skullpatrol):

np :)

OpenStudy (skullpatrol):

@myko Your "proof" assumes that "IF a/0 =b then..." but using 0 in the denominator is strictly not allowed by the definition of division itself.

OpenStudy (anonymous):

my proof says "In case a/0=b then ...". In other words, " supose a/0=b, then ..." . And you get to imposible situation, so the initial asumption have to be wrong

OpenStudy (anonymous):

@skullpatrol

OpenStudy (skullpatrol):

But this case is not allowed by the very definition of division.

OpenStudy (anonymous):

that's what I prooved, because división is defind as a*(1/b) = a/b

OpenStudy (skullpatrol):

Division is defined as a/b= a*(1/b) ; b=/=0 note: b=/=0 is IN the definition.

OpenStudy (skullpatrol):

That is why using 0 for b is UNdefined by the definition.

OpenStudy (anonymous):

read this http://en.wikipedia.org/wiki/Division_by_zero

OpenStudy (skullpatrol):

I have read that.

OpenStudy (anonymous):

so, if you agree with that, we are set, :)

OpenStudy (skullpatrol):

I'm not saying I disagree. I'm saying it is a matter of definition.